U.S. Bureau of the Census data reveal the largest town in the State of OH has a population of one million families. The data show that acreage owned per family is normally distributed with a mean of 10.0 acres and a standard deviation of 1.5 acres. Calculate the following:
a. The percentage of families that own at least 8.1 acres.
b. The number of families that own at least 8.1 acres.
c. The percentage of families that own between 7.0 and 8.1 acres.
d. The minimum acreage owned by the 5% of families owning the largest amount of acreage.
e. If one family from the State of OH is selected at random, what is the probability that family owns more than 8.1 acres or less than 7.0 acres?

Answer by ewatrrr(24785)

 
Hi,
mean of 10.0 acres and a standard deviation of 1.5 acres.
using Excel function: NORMSDIST
a. percentage of families that own at least 8.1 acres: NORMSDIST(-.2667)· 100 = 10.263%
b. The number of families that own at least 8.1 acres: .10263(1M) = 102,630
c. The percentage of families that own between 7.0 and 8.1 acres.
(NORMSDIST(-.2667) - NORMSDIST(-2.00))100 = (.10263 -.02275)100 = 7.988%
d. The minimum acreage owned by the 5% of families owning the largest amount of acreage.
using Excel function: NORMSINV(0.05), z = -1.6449
-1.6449 = (X-10)/1.5
7.5327 Acres = X
e. If one family from the State of OH is selected at random, what is the probability that family owns more than 8.1 acres or less than 7.0 acres?
P= (1-.1026) + .0228 = .9202